# Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond

### v1

@@ 4.1 Preference Vector-Based Methods
Let ${\bm{R}}$ represent the gradients of active constraints, i.e, ${\bm{R}}=\big{\{}\nabla r^{(j)}({\bm{\theta}})\big{\}}_{j\in{\mathbb{I}}({\bm{\theta}})}$, with ${\mathbb{I}}({\bm{\theta}})=\{j\in[n]\mid r^{(j)}({\bm{\theta}})\geq-\epsilon\}$ as the active constraint indices. Using a method akin to MGDA [6], PMTL derives a descent direction for problem (28) that optimizes all $m$ objectives while keeping ${\bm{f}}({\bm{\theta}})$ within the sector closer to the preference vector ${\bm{\alpha}}^{(i)}$. The descent direction is:
@@
where ${\bm{G}}=[\nabla f_{1}({\bm{\theta}}),\cdots,\nabla f_{m}({\bm{\theta}})]$ is the gradient matrix. ${\bm{\lambda}}$ and ${\bm{\beta}}$ are coefficients obtained by solving:
@@
A limitation of PMTL is its weak constraint that only requires the objective vector to lie within a sector, which lacks precise control over the Pareto position.
Exact Pareto Optimal search (EPO) [29, 30] addresses this by ensuring Pareto optimality and aligning the objective vector exactly with a given preference vector ${\bm{\alpha}}$. EPO achieves this using a uniformity function, $u_{{\bm{\alpha}}}({\bm{f}}({\bm{\theta}}))=\text{KL}(\hat{{\bm{f}}}({\bm{\theta}})\parallel\nicefrac{{\bm{1}}}{{m}})$, where: $\hat{f}_{i}({\bm{\theta}})=\frac{\alpha_{i}f_{i}({\bm{\theta}})}{\sum_{j=1}^{m}\alpha_{j}f_{j}({\bm{\theta}})}.$ Minimizing this function aligns ${\bm{f}}({\bm{\theta}})$ with ${\bm{\alpha}}$, achieving precise alignment and Pareto optimality. Let ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ and ${\bm{c}}_{i}$ denote the $i$-th column of ${\bm{C}}$. Denote ${\bm{a}}=[a_{1},\cdots,a_{m}]$, where $a_{i}=\alpha_{i}(\log(m\hat{f}_{i}({\bm{\theta}}))-u_{\bm{\alpha}}(f({\bm{\theta}})))$. At each iteration, EPO classifies objective indices into three sets: ${\mathbb{J}}=\{j\mid{\bm{a}}^{\top}{\bm{c}}_{j}>0\}$ are the indices decreasing uniformity, $\overline{{\mathbb{J}}}=\{j\mid{\bm{a}}^{\top}{\bm{c}}_{j}\leq 0\}$ are the indices increasing uniformity, and ${\mathbb{J}}^{*}$ are the indices with the maximum $\alpha_{j}f_{j}({\bm{\theta}})$. Based on these sets, it determines ${\bm{\lambda}}$ for the common descent direction ${\bm{d}}={\bm{G}}{\bm{\lambda}}$ by solving:
${\bm{\lambda}}=\operatorname*{arg\,max}_{{\bm{\lambda}}\in\Delta_{m-1}}{\bm{\lambda}}^{\top}{\bm{C}}({\bm{a}}\mathbbm{1}_{u_{{\bm{\alpha}}}}+\bm{1}(1-\mathbbm{1}_{u_{{\bm{\alpha}}}})),$
@@
Such updating direction can balance objectives while reducing non-uniformity, resulting in solutions that exactly match each preference vector. However, EPO has three drawbacks: (1) objective vectors must be non-negative, (2) unnecessary complexity from dividing sets into three subsets, and (3) computational inefficiency due to repeated Jacobian calculations and solving linear programming problem (i.e., problem (33)).
To address these limitations, Weighted Chebyshev MGDA (WC-MGDA) [31] considers the dual form of the Tchebycheff scalarization (problem (4)) of a linear programming problem. Optimizing the Tchebycheff function addresses the previously mentioned drawbacks of EPO. However, its convergence rate is relatively slow. Two concurrent approaches have been proposed to improve the Tchebycheff scalarization: (1) the smooth Tchebycheff scalarization (problem (10)) [18] that smooths the original non-smooth $\max$ operation, and (2) the preference-based MGDA (PMGDA) [32] that requires only a negative inner product of the updating direction with the ”exact” constraint gradient, rather than the gradient itself.
@@
The methods discussed above rely on a fixed set of preference vectors. However, in real-world applications, where the shape of the Pareto front is unknown, prefixed preference vectors do not always result in well-distributed solutions. To overcome this limitation, GMOOAR [34] formulates the problem as a bi-level optimization problem. In the upper level, preference vectors are optimized to maximize either the hypervolume or the uniformity of the solutions. In the lower level, solutions are optimized based on these preference vectors. This approach enables the algorithm to dynamically adjust the preference vectors based on the given optimization problem, ensuring the desired distribution of solutions is achieved. UMOD [35] introduces an approach that aims to maximize the minimum distance between solutions:
$\max\left(\min_{1\leq i<j\leq n}\rho({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))\right),~{}~{}\text{s.t.}\;{\bm{\theta}}^{(i)},{\bm{\theta}}^{(j)}\in\mathrm{PS},$ (34)
@@
where $\rho(\cdot,\cdot)$ denotes the Euclidean distance. It achieves desirable solution distributions with a theoretical guarantee: (1) for bi-objective problems with a connected, compact PF, the optimal solution includes the PF endpoints, with equal distances between neighboring objective vectors; (2) as the number of solutions increases, objective vectors asymptotically distribute on the PF.
4.2 Preference Vector-Free Methods
@@
Another line of research does not rely on a preference vector set. Instead, they directly optimize for a diverse set of Pareto optimal solutions.
@@
As HV evaluates solution sets based on both convergence and diversity, it is commonly used for optimizing solution sets [76, 77, 78]. Let $\{{\bm{\theta}}^{(1)},\cdots,{\bm{\theta}}^{(n)}\}$ be a set of solutions and $\{{\bm{f}}({\bm{\theta}}^{(1)}),\cdots,{\bm{f}}({\bm{\theta}}^{(n)})\}$ be their corresponding objective vectors. Using the chain rule, the hypervolume gradient [36, 37, 38] of the $i$-th solution ${\bm{\theta}}^{(i)}$ can be expressed as:
${\bm{d}}^{(i)}=\sum_{k=1}^{m}\underbrace{\frac{\partial\mathrm{HV}_{\bm{r}}(\{{\bm{f}}({\bm{\theta}}^{(1)}),\cdots,{\bm{f}}({\bm{\theta}}^{(n)})\})}{\partial f_{k}({\bm{\theta}}^{(i)})}}_{a_{i}:1\times 1}\underbrace{\frac{\partial f_{k}({\bm{\theta}}^{(i)})}{\partial{\bm{\theta}}^{(i)}}}_{{\bm{B}}:1\times d}.$ (35)
@@
Once ${\bm{d}}^{(i)}$ is calculated, solutions ${\bm{\theta}}^{(i)}$’s are updated by gradient ascent, ${\bm{\theta}}^{(i)}\leftarrow{\bm{\theta}}^{(i)}-\eta{\bm{d}}^{(i)}$, where $\eta$ is a learning rate. Term ${\bm{B}}$ can be easily estimated via backpropagation. The main challenge lies in estimating the first term $a_{i}$. Following Emmerich et al. [38], index sets are categorized based on differentiability: ${\mathbb{Z}}$ includes subvectors with zero partial gradients, ${\mathbb{U}}$ contains subvectors with undefined or indeterminate gradients, and ${\mathbb{P}}$ has subvectors with positive gradients. This classification leads to many zero-gradient terms, improving efficiency. Then, non-zero terms are computed using a fast dimension-sweeping algorithm [79, 80]. This method applies to bi-, tri-, and four-objective problems with time complexities of $\Theta(mn+n\log n)$ for $m=2$ or $3$ and $\Theta(mn+n^{2})$ for $m=4$, where $m$ is the number of objectives and $n$ is the number of solutions. To further enhance convergence, the Newton-hypervolume method [81] incorporates second-order information for faster updates. Deist et al.  [36] apply the hypervolume gradient methods for deep learning tasks such as medical image classification with more than thousands of parameters as decision variables.
@@
MOO-SVGD [39] proposes a different approach inspired by Stein Variational Gradient Descent (SVGD) [82] to maintain diversity while enhancing convergence. The update direction for the $i$-th solution ${\bm{\theta}}^{(i)}$ is given by:
${\bm{d}}^{(i)}=\underbrace{-\sum_{j=1}^{n}{\bm{g}}^{*}({\bm{\theta}}^{(j)})\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))}_{\mathrm{term1}}+$
$\qquad\qquad\qquad\qquad\mu\underbrace{\nabla_{{\bm{\theta}}^{(i)}}\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))}_{\mathrm{term2}},$ (36)
@@
where ${\bm{g}}^{*}({\bm{\theta}}^{(j)})=\sum_{i=1}^{m}\lambda_{i}^{*}\nabla f_{i}({\bm{\theta}}^{(j)}),$ and $\lambda_{i}^{*}$ are the objective weights obtained by MGDA [6] (i.e., problem (13)). The function $\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))$ represents a kernel function that operates on the objective vectors ${\bm{f}}({\bm{\theta}}^{(i)})$ and ${\bm{f}}({\bm{\theta}}^{(j)})$. A common choice for this kernel is the radial basis function (RBF) kernel. In this equation, term1 guides a solution to the PS, while term2 pushes the solutions away from each other to enhance solution diversity.
@@
The methods discussed above focus on scenarios where the number of solutions $n$ typically exceeds the number of objectives $m$. Recent works [40, 42, 83, 41] explore cases with more objectives than solutions (i.e., $m>n$). Few for Many (F4M) [40, 83] identifies solutions using a Tchebycheff method to minimize the maximal value of each objective, formulated as:
$\min_{{\bm{\theta}}^{1},\cdots,{\bm{\theta}}^{n}}\max_{i\in[m]}\left(\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\cdots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}})\right).$ (37)
Sum of Minimum (SoM) [41] minimizes the sum of the minimum values for each objective across solutions, formulated as:
$\min_{{\bm{\theta}}^{1},\cdots,{\bm{\theta}}^{n}}\sum_{i=1}^{m}\left(\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\cdots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}})\right).$ (38)
The difference between F4M and SoM is that the former uses a ”max” aggregation in the outer loop, while the latter uses a ”sum” aggregation.
MosT [42] addresses this problem by finding a uniform Pareto set through a bilevel optimization framework. The outer loop uses a weighted MGDA to find Pareto optimal solutions, while the inner loop optimizes weights via an optimal transport problem for uniform distribution.
@@ 4.3 Discussions
Preference vector-based methods decompose the problem into multiple independent problems based on preference vectors. However, selecting suitable preference vectors can be challenging in real-world scenarios. On the other hand, preference vector-free methods avoid this challenge by directly optimize a set of solutions. While they produce a set of Pareto optimal solutions with interdependent information between solutions, they typically demand more memory and are more computationally intensive.
@@ 5 Finding an infinite set of solutions
Ma et al. [84] initially investigated deriving an infinite set of solutions by using a first-order expansion around discrete Pareto optimal solutions to approximate the PS. However, this method has several drawbacks: (1) the approximation error grows when the solutions are widely spaced, (2) first-order approximations perform poorly in high-dimensional objective spaces, and (3) the approach requires solving a linear programming problem for updates, which can be computationally challenging.
To address these issues, many methods have been introduced that leverage neural networks to learn mappings from user preferences to solutions, enabling the capture of the entire Pareto set. These methods rely on designing efficient network architectures and implementing effective training strategies.
In Sections 5.1.1, 5.1.2, and 5.1.3, we introduce various network structures designed to capture the entire Pareto set through preference-based mappings. Specifically, current algorithms can be categorized into three main types based on their network structures: Hypernetwork-Based, Preference-Conditioned Network-Based, and Model Combination-Based. An illustration of these methods is shown in Figure 4. Finally, in Section 5.2, we discuss training strategies for these network structures.

### v2

@@ 4.1 Methods Based on Preference Vectors
Let ${\bm{R}}$ be the matrix containing gradients of the active constraints, i.e, ${\bm{R}}=\big{[}\nabla r^{(j)}({\bm{\theta}})\big{]}_{j\in{\mathbb{I}}({\bm{\theta}})}$, with ${\mathbb{I}}({\bm{\theta}})=\{j\in[n]\mid r^{(j)}({\bm{\theta}})\geq-\epsilon\}$ containing indices of the active constraints. Using a method akin to MGDA [6], PMTL derives a descent direction for problem (25) that optimizes all $m$ objectives while keeping ${\bm{f}}({\bm{\theta}})$ within the sector closer to the preference vector ${\bm{\alpha}}^{(i)}$. The descent direction is:
@@
where ${\bm{G}}=[\nabla f_{1}({\bm{\theta}}),\dots,\nabla f_{m}({\bm{\theta}})]$ is the gradient matrix. ${\bm{\lambda}}$ and ${\bm{\beta}}$ are coefficients obtained by solving:
@@
A limitation of PMTL is its weak constraint that only requires the objective vector to lie within sectors, which lacks precise control over the position of Pareto solutions.
Exact Pareto Optimal search (EPO) [37, 38] is designed to locate Pareto solutions which are aligned the objective vector exactly with given preference vectors. EPO achieves this using a uniformity function, $u_{{\bm{\alpha}}}({\bm{f}}({\bm{\theta}}))=\text{KL}(\hat{{\bm{f}}}({\bm{\theta}})\parallel\nicefrac{{\bm{1}}}{{m}})$, where: $\hat{f}_{i}({\bm{\theta}})=\frac{\alpha_{i}f_{i}({\bm{\theta}})}{\sum_{j=1}^{m}\alpha_{j}f_{j}({\bm{\theta}})}.$ Minimizing this function aligns ${\bm{f}}({\bm{\theta}})$ with ${\bm{\alpha}}$, achieving precise alignment and Pareto optimality. Let ${\bm{C}}={\bm{G}}^{\top}{\bm{G}}$ and ${\bm{c}}_{i}$ be the $i$-th column of ${\bm{C}}$. Let ${\bm{a}}=[a_{1},\dots,a_{m}]^{\top}$, where $a_{i}=\alpha_{i}(\log(m\hat{f}_{i}({\bm{\theta}}))-u_{\bm{\alpha}}(f({\bm{\theta}})))$. At each iteration, EPO classifies objective indices into three sets: ${\mathbb{J}}=\{j\mid{\bm{a}}^{\top}{\bm{c}}_{j}>0\}$ are the indices decreasing uniformity, $\overline{{\mathbb{J}}}=\{j\mid{\bm{a}}^{\top}{\bm{c}}_{j}\leq 0\}$ are the indices increasing uniformity, and ${\mathbb{J}}^{*}$ are the indices with the maximum $\alpha_{j}f_{j}({\bm{\theta}})$. Based on these sets, it determines ${\bm{\lambda}}$ for the common descent direction ${\bm{d}}={\bm{G}}{\bm{\lambda}}$ by solving:
${\bm{\lambda}}=$ $\operatorname*{arg\,max}_{{\bm{\lambda}}\in\Delta_{m-1}}~{}{\bm{\lambda}}^{\top}{\bm{C}}({\bm{a}}\mathbbm{1}_{u_{{\bm{\alpha}}}}+\bm{1}(1-\mathbbm{1}_{u_{{\bm{\alpha}}}})),$
@@
This updating direction can balance objectives while reducing non-uniformity, resulting in solutions that exactly match each preference vector. However, EPO has three drawbacks: (1) objective vectors must be non-negative, (2) unnecessary complexity from dividing sets into three subsets, and (3) computational inefficiency due to repeated Jacobian calculations and solving linear programming problem (i.e., problem (30)).
To address these limitations, Weighted Chebyshev MGDA (WC-MGDA) [39] considers the dual form of Tchebycheff scalarization (problem (4)) of a linear programming problem. Optimizing Tchebycheff function addresses the previously mentioned drawbacks of EPO. However, its convergence rate is relatively slow. Two concurrent approaches have been proposed to improve the Tchebycheff scalarization: (1) smooth Tchebycheff scalarization (problem (10)) [18] that smooths the original non-smooth $\max$ operation, and (2) the Preference-based MGDA (PMGDA) [40] that requires only a negative inner product of the updating direction with the ”exact” constraint gradient.
@@
The methods discussed above rely on a fixed set of preference vectors. However, in real-world applications where the shape of the Pareto front is unknown, a predefined set of preference vectors may not always result in well-distributed solutions. To overcome this limitation, GMOOAR [42] formulates the problem as a bi-level optimization problem. In the upper level, preference vectors are optimized to maximize either the hypervolume or uniformity of the solutions. In the lower level, solutions are optimized based on these preference vectors. This approach enables the algorithm to dynamically adjust preference vectors based on the given optimization problem, ensuring the desired solution distribution. UMOD [43] introduces an approach that aims to maximize the minimum distance between solutions:
$\max\min_{1\leq i<j\leq n}\rho({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)})),~{}~{}\text{s.t.}\;{\bm{\theta}}^{(i)},{\bm{\theta}}^{(j)}\in\mathrm{PS},$ (31)
@@
where $\rho(\cdot,\cdot)$ denotes the Euclidean distance. UMOD achieves desirable solution distributions with theoretical guarantees: (1) For bi-objective problems with a connected, compact PF, the optimal solution includes the PF endpoints, with equal distances between neighboring objective vectors; (2) As the number of solutions increases, the objective vectors asymptotically distribute on the PF.
4.2 Methods without Using Preference Vectors
@@
Another line of research, directly optimize for a diverse set of Pareto-optimal solutions, does not rely on a given set of preference vectors.
@@
As HV evaluates solution sets based on both convergence and diversity, it is commonly used in traditional evolutionary MOO [86]. In the context of gradient-based MOO, gradient-based hypervolume maximization algorithms (GradHV) [47, 44, 45, 46] have been developed. Let $\{{\bm{\theta}}^{(1)},\dots,{\bm{\theta}}^{(n)}\}$ be a set of solutions and $\{{\bm{f}}({\bm{\theta}}^{(1)}),\dots,{\bm{f}}({\bm{\theta}}^{(n)})\}$ be their corresponding objective vectors. These algorithm first calculate the hypervolume gradient of the $i$-th solution ${\bm{\theta}}^{(i)}$ as follows:
${\bm{d}}^{(i)}=\sum_{k=1}^{m}\underbrace{\frac{\partial\mathrm{HV}_{\bm{r}}(\{{\bm{f}}({\bm{\theta}}^{(1)}),\dots,{\bm{f}}({\bm{\theta}}^{(n)})\})}{\partial f_{k}({\bm{\theta}}^{(i)})}}_{a_{i}:1\times 1}\underbrace{\frac{\partial f_{k}({\bm{\theta}}^{(i)})}{\partial{\bm{\theta}}^{(i)}}}_{{\bm{B}}:1\times d}.$ (32)
@@
Once ${\bm{d}}^{(i)}$ is calculated, solutions ${\bm{\theta}}^{(i)}$’s are updated by gradient ascent, ${\bm{\theta}}^{(i)}\leftarrow{\bm{\theta}}^{(i)}+\eta{\bm{d}}^{(i)}$, where $\eta$ is a learning rate. The term ${\bm{B}}$ in Equation (32) can be easily estimated via backpropagation. The main challenge lies in estimating the first term $a_{i}$. Following Emmerich et al. [46], index sets are categorized based on differentiability: ${\mathbb{Z}}$ includes subvectors with zero partial gradients, ${\mathbb{U}}$ contains subvectors with undefined or indeterminate gradients, and ${\mathbb{P}}$ has subvectors with positive gradients. This classification leads to many zero-gradient terms, improving efficiency. Then, non-zero terms are computed using a fast dimension-sweeping algorithm [87, 88]. This method applies to bi-, tri-, and four-objective problems with time complexities of $\Theta(mn+n\log n)$ for $m=2$ or $3$ and $\Theta(mn+n^{2})$ for $m=4$, where $m$ is the number of objectives and $n$ is the number of solutions. To further enhance convergence, the Newton-hypervolume method [89] incorporates second-order information for faster updates. Deist et al.  [44] apply the hypervolume gradient methods for deep learning tasks such as medical image classification with more than thousands of parameters.
@@
Inspired by Stein Variational Gradient Descent (SVGD) [90], MOO-SVGD [48] proposes a different approach to maintain diversity while enhancing convergence. The update direction for the $i$-th solution ${\bm{\theta}}^{(i)}$ is given by:
${\bm{d}}^{(i)}=$ $-\sum_{j=1}^{n}{\bm{g}}^{*}({\bm{\theta}}^{(j)})\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)}))$
$\qquad\qquad+\mu\nabla_{{\bm{\theta}}^{(i)}}\kappa({\bm{f}}({\bm{\theta}}^{(i)}),{\bm{f}}({\bm{\theta}}^{(j)})).$ (33)
@@
where ${\bm{g}}^{*}({\bm{\theta}}^{(j)})=\sum_{i=1}^{m}\lambda_{i}^{*}\nabla f_{i}({\bm{\theta}}^{(j)}),$ and $\{\lambda_{i}^{*}\}_{i=1}^{m}$ are the objective weights obtained by MGDA [6] (i.e., problem (13)). $\kappa(\cdot,\cdot)$ represents a kernel function, often chosen as the radial basis function (RBF) kernel. The first term in (33) guides the solution towards the PS, while the second term pushes the solutions away from each other to enhance the diversity.
@@
The methods discussed above focus on scenarios where the number of solutions $n$ exceeds the number of objectives $m$. Recent works [49, 51, 91, 50] explore the case with more objectives than solutions (i.e., $m>n$). These approaches assign each objective to a solution ${\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}$ that achieves the smallest objective value for that particular objective. They then minimize either the maximum or the sum of these smallest objective values. Few for Many (F4M) [49, 91] uses the maximum aggregation, formulated as:
$\min_{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}}\max_{i\in[m]}\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}}).$ (34)
Sum of Minimum (SoM) [50] uses the sum aggregation, formulated as:
$\min_{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}}\sum_{i=1}^{m}\left(\min_{{\bm{\theta}}\in\{{\bm{\theta}}^{1},\dots,{\bm{\theta}}^{n}\}}f_{i}({\bm{\theta}})\right).$ (35)
MosT [51] frames the problem as a bilevel optimization task, with the outer loop using MGDA to optimize solutions and the inner loop solving an optimal transport problem to assign objectives to solutions.
@@ 4.3 Discussions
This section reviews two primary approaches for obtaining multiple Pareto optimal solutions.
Preference vector-based methods decompose an MOO problem into multiple independent subproblems based on predefined preference vectors. These subproblems can be solved independently, which enhances memory efficiency and enables parallel computation. However, in real-world applications, determining suitable preference vectors can be challenging due to the complex and often unpredictable nature of the Pareto front. Furthermore, since the subproblems are solved independently, they cannot benefit from shared information between one another.
On the other hand, methods not using preference vectors bypass the challenge of selecting preference vectors by directly optimizing a set of solutions. These methods generate a set of Pareto-optimal solutions that inherently share interdependent information. However, this approach typically requires higher memory resources and incurs greater computational costs, making it less efficient in resource-constrained scenarios.
@@ 5 Finding an Infinite Set of Solutions
Ma et al. [92] initially investigated deriving an infinite set of solutions by approximating the Pareto set with first-order expansion around discrete Pareto-optimal solutions. However, this method has several drawbacks: (1) The approximation error grows when the solutions are widely spaced, (2) First-order approximations perform poorly in high-dimensional objective spaces, and (3) The approach requires solving a linear programming problem for updates, which can be computationally challenging.
To address these issues, many methods have been introduced that leverage neural networks to learn mappings from user preferences to solutions directly, enabling the capture of the entire PS. These methods rely on designing efficient network architectures and implementing effective training strategies.
In Section 5.1, we introduce various network structures designed to capture the entire PS through preference-based mappings. Specifically, current algorithms mainly use three types of network structures: (1) Hypernetwork in Section 5.1.1, (2) Preference-Conditioned Network in Section 5.1.2, and (3) Model Combination in Section 5.1.3. An illustration of these methods is shown in Figure 5. Section 5.2 discusses training strategies for these network structures.